### v1

# E-Gen: Leveraging E-Graphs to Improve Continuous Representations of Symbolic Expressions

Abstract
As vector representations have been pivotal in advancing natural language processing (NLP), some prior research has concentrated on creating embedding techniques for mathematical expressions by leveraging mathematically equivalent expressions. While effective, these methods are limited by the training data. In this work, we propose augmenting prior algorithms with larger synthetic dataset, using a novel e-graph-based generation scheme. This new mathematical dataset generation scheme, E-Gen, improves upon prior dataset-generation schemes that are limited in size and operator types. We use this dataset to compare embedding models trained with two methods: (1) training the model to generate mathematically equivalent expressions, and (2) training the model using contrastive learning to group mathematically equivalent expressions explicitly. We evaluate the embeddings generated by these methods against prior work on both in-distribution and out-of-distribution language processing tasks. Finally, we compare the performance of our embedding scheme against state-of-the-art large language models and demonstrate that embedding-based language processing methods perform better than LLMs on several tasks, demonstrating the necessity of optimizing embedding methods for the mathematical data modality.
E-Gen: Leveraging E-Graphs to Improve Continuous Representations of Symbolic Expressions
Hongbo Zhengthanks: Equal Contribution.  Suyuan Wang11footnotemark: 1  Neeraj Gangwar  Nickvash Kani University of Illinois at Urbana-Champaign {hongboz2,suyuan2,gangwar2,kani}@illinois.edu
1 Introduction
While large language models (LLMs) (Devlin, 2018; Radford, 2018; Radford et al., 2019; Rae et al., 2021; Chowdhery et al., 2023; Achiam et al., 2023) have demonstrated effectiveness in processing natural language, these methods still perform suboptimally with math-based content which plays an important role across numerous domains (Zanibbi et al., 2024; Rohatgi et al., 2019). For instance, even a state-of-the-art LLM like GPT-4V only reaches 49% accuracy in MathVista (Lu et al., 2023), a comprehensive benchmark to evaluate the mathematical reasoning capabilities of LLMs. In another instance, as tested in Frieder et al. (2024), GPT-4’s performance in computing integration is dominated by specialized embedding-based models (Lample and Charton, 2019; Lample et al., 2022). Therefore, effective approaches to process semantically rich mathematical notation are necessary.
One approach is to compute a semantic representation of mathematical content based on textual context and treat mathematical expressions as a sentence without learning mathematical semantics (Krstovski and Blei, 2018). This has been used in various mathematical language processing (MLP) tasks, such as mathematical information retrieval (Topić et al., 2013; Mansouri et al., 2022a), identifier definition extraction (Pagael and Schubotz, 2014; Hamel et al., 2022; Zou et al., 2024), mathematical reasoning (Geva et al., 2020; Nye et al., 2021) and theorem proving (Wang et al., 2020; Wu et al., 2022). But this approach has two limitations: (1) textual descriptions are lacking in some math content like textbooks or mathematical derivations in academic publications, and (2) don’t really explore relations between symbolic operators from a mathematical perspective.
To address these problems, recent studies have focused on equivalence and mathematical manipulation between expressions to derive the meaning of mathematical expressions independent of context. However, these approaches remain limited, primarily due to the quality of available data. Allamanis et al. (2017) introduces EqNet, a TreeNN-based (Socher et al., 2013) model to group equivalent expressions together, but their datasets are limited to arithmetic and boolean expressions. SemEmb (Gangwar and Kani, 2023) computes vector-based semantic representations by learning to generate mathematically equivalent expressions but is significantly limited by the SymPy-generated dataset (Meurer et al., 2017). SymPy functions are designed for simplification and directly derive the most simplified expression without intermediate steps, leading to a dataset that lacks enough rewrites per expression. Hence, a more efficient mathematical data generation scheme is required.
(a) Initial e-graph
$(x+8)-8$
(b) After applying rule
$(x+y)-z\rightarrow x+(y-z)$
(c) After applying rule
$x-x\rightarrow 0$
(d) After applying rule
$x+0\rightarrow x$
(e) Grammar created from
saturated e-graph
Figure 1: Illustration of e-graph saturation (1(a) to 1(d)) and grammar creation (1(e)). An e-graph consists of e-classes (dashed boxes) containing equivalent e-nodes (solid boxes). Edges connect e-nodes to their child e-classes. Applying mathematical rules to an e-graph adds new e-nodes and edges (1(b) and 1(c)), or merges e-classes (1(d)). Additions and modifications are emphasized in black. In 1(e), the saturated e-graph is converted into a context-free grammar, where each grammar is defined using e-class id and the e-nodes with their child e-classes.
In this work, we propose E-Gen, a highly scalable and efficient corpus generation scheme based on e-graphs (Willsey et al., 2021). Leveraging a collection of mathematical transformations, E-Gen facilitates the creation of synthetic datasets with large clusters of semantically equivalent expressions. This approach significantly improves upon prior SymPy-based methods, overcoming limitations in the number of rewrites per expression while improving flexibility and scalability. In summary, we highlight the following contributions of E-Gen:
  1. 1)
    We introduce E-Gen, a novel scheme for generating a cluster-based mathematical expression dataset, along with a high-diversity mathematical corpus.
  2. 2)
    We evaluate two types of embedding models based on seq2seq and contrastive learning respectively, showing improved semantic representation performance over prior works in quantitative and qualitative tests.
  3. 3)
    The embedding models are evaluated on two out-of-distribution (OOD) downstream tasks, to demonstrate models’ generalizability and robustness.
  4. 4)
    Finally, we compare our models with GPT-4o across multiple tasks, demonstrating the effectiveness of embedding-based approaches.
2 Related Work
In mathematical language processing (MLP) (Meadows and Freitas, 2022), semantic representation shows strong potential across various problems. A representative application is mathematical information retrieval (MIR) (Kristianto et al., 2016; Zanibbi et al., 2016a; Mansouri et al., 2022c), where expressions or keywords are ranked based on their relevance to a query. Gao et al. (2017) proposes Symbol2Vec, a mathematical symbol representation generation scheme to group LaTeX symbols having similar contexts together. They further extend this approach to a MIR scheme (Formula2Vec) to prove its effectiveness. Krstovski and Blei (2018) converts symbolic layout trees (Zanibbi et al., 2016b) of mathematical expressions into token sequences and generates representation based on word2vec. Mansouri et al. (2019) further improves this approach by combining the symbol layout tree and operator tree representations of an expression. Peng et al. (2021) proposes MathBERT, a pretrained model based on MIR and formula headline generation task (Yuan et al., 2020).
Identifier definition extraction (Kristianto et al., 2014) is another promising application, which aims to align identifiers found in scientific text with their definitions. Popovic et al. (2022) utilizes a transformer-based method to develop an end-to-end joint math entity and relation extraction approach. Jo et al. (2021) provides a mathematical understanding pretrained model by fine-tuning BERT on masked mathematical identifier prediction task. In the theorem proving field, Welleck et al. (2021b, a) generate theorem proof by retrieving relevant references to a given query theorem. However, as discussed in Section 1, MLP studies rely on semantic representation but mostly focus on establishing homomorphism between context and mathematical expressions to construct semantic representation, which is limited by the incapacity to process pure math content and can not really understand mathematical transformations. Even though some recent studies (Allamanis et al., 2017; Meidani et al., 2023) have focused on generating semantic representation based on intrinsic features of mathematical expressions to address this problem, their performance and scope of application are still limited by existing dataset, motivating this work’s demonstration of a more effective mathematical corpus generation scheme.
3 Corpus of Equivalent Expressions
3.1 Corpus Generation
To generate a diverse set of mathematical expressions, we manually design templates containing placeholders for arithmetic operators, functional operators, and numerical values. Expressions are systematically instantiated by replacing the placeholders in templates with all possible operators and values, resulting in approximately 5,000 raw expressions in prefix notation. The templates are carefully designed to cover fundamental arithmetic and functional operators while maximizing the potential for mathematical transformations. Each of the resulting raw expressions is processed by E-Gen, which applies around 800 mathematical rules to generate clusters of semantically equivalent expressions.
The core of our E-Gen is e-graph, an advanced data structure designed to efficiently manipulate collections of terms under a congruence relation. An e-graph is composed of e-classes, each containing a set of equivalent e-nodes. An e-node can be linked to one or more child e-classes, depending on the operator’s arity. From a mathematical perspective, child e-classes represent the arguments of their associated e-node, typically corresponding to a mathematical operator. Consequently, any subgraph originating from an e-node within the same e-class represents equivalent expressions.
Figures 1(a) to 1(d) show the e-graph saturation process. In the initial e-graph (Figure 1(a)), the e-node “$+$” is linked to two child e-classes “$x$” and “$8$”, respectively, and the e-node “--” is linked to a subgraph “$(x+8)$” and a child e-class “$8$”, together forming the raw expression “$(x+8)-8$”. In Figure 1(b), a new e-node “$+$” is added to the top e-class after the associative law is introduced. The two e-nodes, “--” and “$+$” in the top e-class represent subgraphs for expressions “$(x+8)-8$” and “$x+(8-8)$” which are equivalent rewrites generated by the associative law. Similarly, transformations such as “$(x-x)\rightarrow 0$” and “$x+0\rightarrow x$” are embedded into the e-graph in the steps shown in Figures 1(c) and 1(d). The e-graph is iteratively expanded by applying each applicable rule, thereby capturing all possible equivalent expressions.
A context-free grammar is created after e-graph saturation in prefix notation, based on e-classes and the connections of e-nodes in it, as shown in Figure 1(e). E-classes will be represented as variable symbols denoted by "$e0$", "$e1$", "$e2$", "$e4$" in Figure 1(d). Variables and numbers comprise the terminal set in the grammar. The production rules are determined by the edges and take the form:
$[ec_{i}]$ $\rightarrow[var/num]$
$[ec_{i}]$ $\rightarrow[op\;ec_{j}\;ec_{k},\ldots]$
where $ec$ denotes eclasses, $op$ and $var/num$ denotes enodes representing operators and variables/numbers.
[…]
Figure 2: Overview of E-Gen and the cluster-based training framework. Equivalent expressions of each initial expression are generated using an e-graph-based approach, forming clusters of equivalent expressions. The seq2seq model is trained on equivalent expression pairs, while the contrastive learning model is trained on triplets, with each triplet containing a reference expression $\boldsymbol{x}$, an equivalent positive sample $\boldsymbol{x}^{+}$, and a non-equivalent negative sample $\boldsymbol{x}^{-}$. $\boldsymbol{z}$ is the latent space representation of the corresponding input $\boldsymbol{x}$.
3.2 Corpus Analysis
Category Operators
Arithmetic $+$, --, $\times$, $\div$
$\operatorname{pow}$, $\operatorname{abs}$, $\operatorname{sqrt}$, $\frac{d}{dx}$
Logarithmic/Exponential $\ln$, $\exp$ (as $\operatorname{pow}e$)
Trigonometric $\sin$, $\cos$, $\tan$
$\csc$, $\sec$, $\cot$
Inverse Trigonometric $\sin^{-1}$, $\cos^{-1}$, $\tan^{-1}$
$\csc^{-1}$, $\sec^{-1}$, $\cot^{-1}$
Hyperbolic $\sinh$, $\cosh$, $\tanh$
$\operatorname{csch}$, $\operatorname{sech}$, $\coth$
Inverse Hyperbolic $\sinh^{-1}$, $\cosh^{-1}$, $\tanh^{-1}$
$\operatorname{csch}^{-1}$, $\operatorname{sech}^{-1}$, $\coth^{-1}$
Table 1: Operator coverage of E-Gen Corpus.
The new corpus encompasses a comprehensive set of arithmetic and functional operators as shown in Table 1. Table 2 presents partial examples from two clusters of equivalent expressions generated by E-Gen. The corpus is structured into clusters, where each cluster contains numerous mathematically equivalent expressions derived through a series of flexible transformations. These transformations span both simple and complex mathematical relationships, significantly enhancing expression diversity and enabling models to learn underlying mathematical rules more efficiently with fewer raw expressions.
Raw expression: $\tanh(3x-(-4))-6$
$\tanh(3x+4)-6$
$61/\coth(3x+4)-$ 6
$\sinh(3x+4)\operatorname{sech}(3x+4)-6$
$\tanh(3/\csc(\sin^{-1}({x}))+4)-6$
$\sinh(3\cos(\cos^{-1}({x}))+4)/\cosh(3x+4)-6$
Raw expression: $\frac{d}{dx}-2(\ln x/7)^{-2}$
$4/\left(x(\ln(x/7))^{3}\right)$
$-2/\left(x/(\ln(x)-\ln(7))^{-3}\right)\times(-2)$
$-4x^{-1}/\ln(7\times x^{-1})^{3}$
$-4/(x(\ln(7)-\ln(x))^{3})$
$4\cot(\tan^{-1}(x))(\ln(x/7))^{-3}$
Table 2: Examples of equivalent expressions generated with E-Gen. Expressions listed below each raw expression are part of the equivalent rewrites. Additional examples are in Appendix A.2.
As detailed in Table 3, the E-Gen corpus not only supports a wider range of operators, but also achieves a significantly higher average cluster size of 102, compared to an existing corpora Equivalent Expressions Dataset (EED) (Gangwar and Kani, 2023) that is generated using SymPy and limited to a cluster size of 2 (pairs). This substantial increase in cluster size enables models to develop a deeper semantic understanding of the diversity of mathematical transformations.
Statistic E-Gen EED
# Operators 34 24
Average Cluster Size 102 2
Average Sequence Length 15 16
Train Set Size (Pair) $\sim$ 55M $\sim$ 4.66M
Train Set Size (Triplet) $\sim$ 50M -
Test Set Size 8,077 5,000
Table 3: Dataset Statistics. Corpus comparison between E-Gen and EED which denotes Equivalent Expressions Dataset from prior work (Gangwar and Kani, 2023).
Training Data.
The training dataset is provided in two formats: expression pairs and triplets, corresponding to the two training methodologies described in Section 4. For the seq2seq model, equivalent expression pairs are generated by permuting expressions within each cluster. For contrastive learning, expression triplets are constructed by treating expressions within the same cluster as positive pairs, while randomly sampling expressions from different clusters to serve as negative examples. This process results in a training set comprising 55 million equivalent expression pairs and 50 million expression triplets as shown in Table 3. For the validation set, a subset of clusters is randomly sampled from the corpus, yielding a total of 8,077 expressions. Following Lample and Charton (2019) and Gangwar and Kani (2023), we use the prefix notation to encode the expressions.
4 Methodology
As illustrated in Figure 2, we employ two well-established approaches for learning representations of mathematical expressions: sequence-to-sequence (seq2seq) (Sutskever, 2014; Cho, 2014) equivalent expression generation and contrastive learning (CL) (Wu et al., 2018; Chen et al., 2020), both leveraging the transformer architecture (Vaswani, 2017). These methods aim to capture the underlying semantic relationships between mathematical expressions by generating meaningful embeddings in a high-dimensional latent space.
4.1 seq2seq Expressions Generation
Following prior work (Gangwar and Kani, 2023), we employ the seq2seq framework to learn representations of mathematical expressions by training the model to generate mathematically equivalent expressions. Specifically, given a pair of equivalent expressions, the model takes one as input and is tasked with predicting the other as output. The encoder, during this process, learns to map the input expressions into a latent space where semantically equivalent expressions are clustered together.
4.2 Contrastive Learning
The other promising approach to learning mathematical expression representations is contrastive learning, a technique that has gained significant traction in the domain of representation learning. The primary objective is to learn a latent space where semantically equivalent expressions are embedded closer together, while semantically distinct expressions are pushed apart. This is achieved using a contrastive loss function, such as InfoNCE (Oord et al., 2018) or SimCSE (Gao et al., 2021).
In this manuscript, we use a variation of the InfoNCE loss formulated as:
$\mathcal{L}(f)=\mathbb{E}\left[-\ln\frac{e^{f^{\intercal}(\boldsymbol{x})f(\boldsymbol{x}^{+})}/\tau}{e^{f^{\intercal}(\boldsymbol{x})f(\boldsymbol{x}^{+})}/\tau+e^{f^{\intercal}(\boldsymbol{x})f(\boldsymbol{x}^{-})}/\tau}\right]$ (1)
where $f$ is a transformer encoder $f:\mathcal{X}\rightarrow\mathcal{Z}$ that maps an tokenized input expression $\boldsymbol{x}\in\mathcal{X}$ to its latent representation $\boldsymbol{z}\in\mathcal{Z}$. The terms $\boldsymbol{x}^{+}$ and $\boldsymbol{x}^{-}$ correspond to positive (equivalent) and negative (non-equivalent) samples respectively, with $\tau$ as a temperature hyperparameter.
4.3 Representation Vector
To derive the representation vector from the transformer encoder, an additional step is required to convert the output matrix from the encoder’s last layer $\boldsymbol{X}\in\mathbb{R}^{S\times D_{\text{model}}}$ to a one-dimensional embedding vector $\boldsymbol{x}\in\mathbb{R}^{\textit{D}_{\text{model}}}$ for each expression. Here, $S$ and $D_{\text{model}}$ represent the sequence length of the input and the model dimension. We experiment with two common pooling strategies: average pooling and max pooling over the hidden states of the last encoder layer. Our empirical analysis and prior work (Gangwar and Kani, 2023) indicate that max pooling consistently outperforms average pooling for the seq2seq model. In the case of contrastive learning, we train two separate models—one with average pooling (CL Mean) and one with max pooling (CL Max)—to evaluate the impact of these strategies on embedding quality. In all models, special tokens, such as the start-of-expression and end-of-expression tokens, are excluded from the pooling process.
5 Experiments
5.1 Evaluation Tasks
The models trained on the E-Gen synthetic dataset, are compared against the prior SemEmb model (Gangwar and Kani, 2023) trained on SymPy-generated corpus, on both in-distribution and out-of-distribution tasks, demonstrating the efficacy of the newly presented methods.
K-Means Clustering.
Figure 3: Visualization of representation vectors for 17 different single-operator mathematical expressions and their equivalent forms using our method (left) and SemEmb (right). t-SNE (Van der Maaten and Hinton, 2008) is applied to reduce the dimensionality of the embeddings from 512 to 2.
K-Means clustering (Macqueen, 1967) is employed to evaluate the performance of both seq2seq and contrastive learning models. The test set comprises 8,077 expressions in 279 clusters, with cluster sizes ranging from 20 to 40 expressions. Subsequently, K-Means clustering is applied to group the expressions into clusters. A visualization of clusters is shown in Figure 3. To evaluate clustering performance, the predicted cluster labels from the K-Means algorithm are mapped to the corresponding ground truth labels, with accuracy computed as the proportion of expressions correctly assigned to their respective clusters.
Model Accuracy (%)
seq2seq 96.72
CL Mean 97.61
CL Max 97.30
SemEmb 37.70
Table 4: K-Means clustering accuracy (%) of seq2seq, CL Mean, and CL Max, compared against prior SemEmb model.
As shown in Table 4, both the seq2seq and contrastive learning models, trained on the E-Gen cluster-based dataset, demonstrate strong performance in grouping equivalent expressions while separating non-equivalent ones. Notably, the CL models slightly outperform the seq2seq model in this task because they are trained to explicitly group similar expressions together. In contrast, SemEmb, which is trained on a SymPy-generated corpus, only achieves 37.70% accuracy. The lower performance is attributed to the limited diversity of mathematical transformations in the SymPy-generated corpus which restricts the model’s ability to effectively capture and distinguish small and large syntactic variations among equivalent expressions.
Semantic Understanding Beyond Syntactic Similarity.
To evaluate whether models trained on the E-Gen corpus capture mathematical semantics rather than relying on syntactic similarity, we design an experiment where the model must identify the single semantically equivalent expression among syntactically similar candidates. Specifically, we sample 1,000 query expressions from the test set and, for each query, we introduce one correct answer that is semantically equivalent but syntactically distinct from the query. Additionally, six distractors are generated by making minor syntactic modifications over the query and the correct answer (e.g., altering a single operator or numerical value). These distractors are divided into two groups: three closely resembling the query and three closely resembling the correct answer. The model is tasked with selecting the correct equivalent expression from seven candidates. In the example shown in Table 5, the candidate highlighted in bold represents the correct answer, which is syntactically distinct from the query involving the derivative operation. Candidate 4, 5, 7 are syntactically similar to the correct answer, while candidate 1, 3, 6 shares a similar structure with the query. As shown in Table 6, models trained on the E-Gen corpus significantly outperform SemEmb, demonstrating the models’ ability to distinguish between semantically equivalent but syntactically different expressions, as well as syntactically similar yet semantically non-equivalent expressions.
Query expression: $\frac{d}{dx}2x+\sin(x-4)/6$
$\frac{d}{dx}3x+\sin(x-4)/6$
$\boldsymbol{2+\cos(x-4)/6}$
$\frac{d}{dx}2x+\sin(x+4)/6$
$62+\cos(x/4)/$ 6
$82+\tan(x-4)/$ 8
$\frac{d}{dx}2x+\cos(x-4)/6$
$69+\cos(x-4)/$ 6
Table 5: Examples of semantic understanding beyond syntactic similarity query, where the correct option is written in red. Six distractors are generated by slightly changing the query expression and the correct option to evaluate model’s capacity in distinguishing syntactically similar but semantically different expressions
Model Accuracy (%)
seq2seq 76.41
CL Mean 50.36
CL Max 48.19
SemEmb 28.62
Table 6: Semantic understanding beyond syntactic similarity accuracy (%) of seq2seq, CL Mean, and CL Max, compared against prior SemEmb model.
$\boldsymbol{x}_{1}$ $\boldsymbol{y}_{1}$ $\boldsymbol{x}_{2}$ $\boldsymbol{\hat{y}}_{2}$ (seq2seq) $\boldsymbol{\hat{y}}_{2}$ (SemEmb) $\boldsymbol{\hat{y}}_{2}$ (GPT-4o) $\boldsymbol{y}_{gt}$
$1$ $\sin(x)$ $-\sin(-x)$ $\cos(x)$ $\cos(-x)$ ${\color[rgb]{1,0,0}-\tan(-x)}$ $\cos(-x)$ $\cos(-x)$
$2$ $\cos(x)$ $\sec(x)$ $\tanh(x)$ $\coth(x)$ $\coth(x)$ ${\color[rgb]{1,0,0}\operatorname{sech}(x)}$ $\coth(x)$
$3$ $\sinh^{-1}(x)$ $\operatorname{csch}^{-1}(1/x)$ $\tanh^{-1}(x)$ $\coth^{-1}(1/x)$ ${\color[rgb]{1,0,0}\coth^{-1}(x)}$ $\coth^{-1}(1/x)$
$4$ $\tan(x)$ $\tan(x+\pi)$ $\csc(x)$ $\csc(x+2\pi)$ ${\color[rgb]{1,0,0}\cot(x+\pi)}$ $\csc(x+2\pi)$ $\csc(x+2\pi)$
$5$ $\sin(x)$ $\cos(x-\pi/2)$ $\sec(x)$ $\csc(x+\pi/2)$ ${\color[rgb]{1,0,0}\sec(x-\pi/2)}$ ${\color[rgb]{1,0,0}\csc(x-\pi/2)}$ $\csc(x+\pi/2)$
$6$ $x$ $\ln{x}$ $\coth(x)$ $\ln{\coth(x)}$ ${\color[rgb]{1,0,0}\coth{\ln(x)}}$ ${\color[rgb]{1,0,0}\coth^{-1}(x)}$ $\ln{\coth(x)}$
$7$ $x$ $\ln{x}$ $\cos^{-1}(x)$ ${\color[rgb]{1,0,0}\cos^{-1}(\ln x)}$ ${\color[rgb]{1,0,0}\cos^{-1}({\ln x})}$ $\ln{\cos^{-1}(x)}$ $\ln{\cos^{-1}(x)}$
$8$ $x$ $\csc^{-1}(x)$ $\operatorname{csch}(x)$ $\csc^{-1}(\csc(x))$ ${\color[rgb]{1,0,0}\operatorname{csch}^{-1}(x)}$ $\csc^{-1}(\csc(x))$
$9$ $x$ $x+1$ $\tan^{-1}(x)$ ${\color[rgb]{1,0,0}\tan^{-1}(x)\times 1}$ $\tan^{-1}(x)+1$ $\tan^{-1}(x)+1$ $\tan^{-1}(x)+1$
$10$ $x$ $x^{3}$ $\sinh^{-1}(x)$ $\sinh^{-3}(x)$ $\sinh^{-3}(x)$ $\sinh^{-3}(x)$ $\sinh^{-3}(x)$
Table 7: Example comparison of embedding algebra predictions between the seq2seq, SemEmb, and GPT-4o. The model’s prediction is denoted as $\hat{\boldsymbol{y}}_{2}$, while $\boldsymbol{y}_{gt}$ represents the ground truth. Incorrect predictions are in red. Additional experimental results are provided in Appendix C.4.
Mistake Detection.
Mistake detection in mathematical derivations is a downstream OOD task of which test set is generated by Sympy to test the robustness and generalizability of the models.
Derivation seq2seq GPT-4o
$-\frac{1}{8}(\sinh(\cosh^{-1}(7-\frac{x}{-8})))^{-1}$
$-\frac{1}{8}(\sinh(\cosh^{-1}(7+\frac{x}{8})))^{-1}$
$-\frac{1}{8}(\sqrt{6+\frac{x}{8}}\sqrt{8+\frac{x}{8}})^{-1}$bigstar
$-\frac{1}{8}(6+\frac{x}{8})^{-\frac{1}{2}}(8+\frac{x}{8})^{-\frac{1}{2}}$
$-\frac{1}{8}(6+\frac{x}{8})^{-\frac{1}{2}}(8+{\color[rgb]{1,0,0}(1-8)}x)^{-\frac{1}{2}}$bigstar $\bigstar$
$-\frac{1}{8}(6+\frac{x}{8})^{-\frac{1}{2}}(8-7x)^{-\frac{1}{2}}$
$5/\csc(\csc^{-1}(1/\ln(\frac{x}{-5})))$
$5/\csc(\csc^{-1}(1/\ln(\frac{1}{-5}x)))$
$5/{\color[rgb]{1,0,0}\sec}(\csc^{-1}((\ln(\frac{1}{-5}x))^{-1}))$bigstar
$5/(1-(\ln(\frac{1}{-5}x))^{2})^{-\frac{1}{2}}$
$5\sqrt{1-(\ln(\frac{1}{-5}x))^{2}}$
Table 8: Example comparison of mistake detection in mathematical derivations between seq2seq and GPT-4o. Errors in the derivations are in red. The $\bigstar$ symbol indicates that the respective model has predicted the step to contain a mistake.
In our work, we first generate a step-by-step mathematical derivation based on expressions in the original test set of the E-Gen corpus. Then, mistakes are introduced to randomly selected steps for each derivation. Models are expected to identify steps with mistakes in the derivation steps. Specifically, transformations between consecutive derivation steps are classified as either “mistake” or “no mistake”. 80% derivations are used to calculate a semantic similarity threshold as the threshold for binary classification [^1] Similarity threshold calculation is shown in Appendix C.2, with the remainder being the test set. A step is flagged as a potential mistake when the cosine similarity between prior steps and the current step is below the threshold. Two examples of mistake detection are shown in Table 8, some steps with minor mistakes often share a similar structure with the preceding steps, such as mistakes steps shown in Table 8, making some incorrect transformations between consecutive steps appear deceptively more similar to correct mathematical transformations that significantly change the expressions’ structure. This poses a significant challenge to models’ ability to understand mathematical semantics.
Model Precision Recall F1
no mistake seq2seq 96.40 94.69 95.54
CL Mean 97.59 92.52 94.99
CL Max 97.78 91.93 89.14
SemEmb 92.92 83.49 87.95
mistake seq2seq 74.68 81.61 77.99
CL Mean 69.33 88.10 77.60
CL Max 67.96 89.14 77.12
SemEmb 44.46 67.50 53.61
Table 9: Mistake detection evaluation results precision, recall, and F1 (%) scores of seq2seq, CL Mean, and CL Max, compared against prior SemEmb model.
The test set has 18,462 derivation steps, including 2,974 steps containing mistakes, with the remainder being error-free. As shown in Table 9, precision, recall, and F1-score are used to evaluate the performance. All three models trained on E-Gen corpus demonstrate strong effectiveness in identifying potential mistakes and better generalizability in OOD task, significantly outperforming the SemEmb approach.
Embedding Algebra.
Embedding algebra is a classic task to evaluate if embeddings capture semantic information of an word/token. Techniques like word2vec (Mikolov, 2013) and GloVe (Pennington et al., 2014) exhibit the ability to perform analogy-based reasoning through algebraic operations on their representation vectors, facilitating solutions to analogies such as “Berlin is to Germany as Paris is to France”. By extending this task to mathematical analogies, we test if models can really understand mathematical transformations or just have a shallow understanding of expressions’ structure. For a given triplet of expressions $\boldsymbol{x}_{1}$, $\boldsymbol{y}_{1}$, and $\boldsymbol{x}_{2}$, we compute:
$f(\boldsymbol{\hat{y}}_{2})=-f(\boldsymbol{x}_{1})+f(\boldsymbol{y}_{1})+f(\boldsymbol{x}_{2})$ (2)
where $f$ denotes a function $f:\mathcal{X}\rightarrow\mathcal{Z}$, which maps an expression $\boldsymbol{x}$ to its representation vector $\boldsymbol{z}$. We then identify the expression whose embedding vector is most similar to $f(\boldsymbol{\hat{y}}_{2})$ in terms of cosine similarity, excluding the original expressions $\boldsymbol{x}_{1}$, $\boldsymbol{y}_{1}$, and $\boldsymbol{x}_{2}$ from consideration.
For this experiment, 584 analogy examples are manually constructed, and all expressions from the E-Gen corpus are used as the search pool, removing those equivalent to $\boldsymbol{x}_{2}$ and $\boldsymbol{y}_{2}$ to ensure uniqueness of answer, while adding any necessary expressions to complete the analogy.
Model Accuracy (%)
seq2seq 70.38
CL Mean 64.73
CL Max 50.34
SemEmb 54.85
GPT-4o 39.60
Table 10: Embedding algebra accuracy (%) of seq2seq, CL Mean, and CL Max, compared against prior SemEmb model and GPT-4o.
As shown in Table 10, the seq2seq model achieves the highest accuracy, indicating its ability to effectively learn underlying mathematical rules and handle basic substitutions. Conversely, SemEmb exhibit low accuracy on this task and have a tendency to imitate the transformation between $\boldsymbol{x}_{1}$ and $\boldsymbol{y}_{1}$, leading to “look-alike” predictions rather than a true understanding of the mathematical rules, such as Test 1, 4, 5. Examples are shown in Table 7.
5.2 Comparison with GPT-4o
To further assess the quality of vector representations of symbolic expressions, a comparative analysis is conducted between the seq2seq model, trained on the E-Gen corpus, and the state-of-the-art large language model GPT-4o (Achiam et al., 2023) on several tasks.
Mistake Detection.
For mistake detection, a small test set is randomly sampled from the mistake detection test set mentioned above, consists of 322 derivation steps, of which 50 contain mistakes and the remainder are error-free. We first provide GPT-4o a derivation example in the prompt, asking it to identify potential mistakes. If GPT-4o fails to find the potential mistake, we teach it where the mistake appears until it can understand the task and pass the test. Once verified, we evaluate GPT-4o using derivation queries from the test set. As shown in Table 11, our work outperforms GPT-4o across all metrics. GPT-4o demonstrates a recall of 66%, suggesting its capability of detecting a notable portion of errors. However, its low precision indicates a tendency to misclassify correct transformations as mistakes, leading to a high rate of false positives.
Model Precision Recall F1
no mistake seq2seq 96.98 94.49 95.72
CL Mean 97.64 91.18 94.30
CL Max 97.78 91.93 94.77
GPT-4o 93.58 90.84 92.19
mistake seq2seq 73.68 84.00 78.50
CL Mean 64.71 88.00 74.58
CL Max 67.96 89.14 77.12
GPT-4o 56.90 66.00 61.11
Table 11: Mistake detection evaluation results precision, recall, and F1 (%) scores of seq2seq, CL Mean, and CL Max compared with GPT-4o.
Table 8 illustrates examples of GPT-4o’s misjudgments, which commonly occur during complex transformations, particularly those involving function operators or relatively intricate arithmetic calculations. For instance, in Example 1, GPT-4o performs suboptimally with evaluating equivalence when expressions are syntactically different (e.g., $\sinh(\cosh^{-1}(x))=\sqrt{x+1}\sqrt{x-1}$). In Example 2, GPT-4o fails to identify the operator substitution in step 3, judging it as a valid derivation.
Embedding Algebra.
We also conduct the embedding algebra task on GPT-4o, which performs worse than both the seq2seq and CL embedding models, achieving only 39.60% accuracy on the test set as shown in Table 10. Table 7 provides examples comparing the performance of the seq2seq model and GPT-4o on embedding algebra tests. Similar to the mistake detection test, an example query is presented to GPT-4o to verify its understanding and confirm that it produces the correct output before proceeding with the test. While the seq2seq model consistently makes accurate predictions by adhering to mathematical rules, GPT-4o shows a partial understanding of certain mathematical properties, such as correctly predicting the periodicity of “$\csc(x)$”. However, it often imitates the structure of “$\boldsymbol{y}_{1}$” without fully considering the underlying mathematical rules. For instance, in Test 5, GPT-4o incorrectly imitates the structure of “$\boldsymbol{y}_{1}$”, subtracts “$\pi/2$” from “$x$” in “$\csc(x)$”, while the correct answer should be “$\csc(x+\pi/2)$” which is mathematically equivalent to “$\csc(x)$”. A similar error occurs in Test 8, where GPT-4o predicts “$\boldsymbol{y}_{2}$” as “$\operatorname{csch}^{-1}(x)$” by imitating “$\boldsymbol{y}_{1}:\csc^{-1}(x)$”, instead of applying the correct mathematical transformation. This comparison highlights the effectiveness of E-Gen corpus in helping models to build up understanding in mathematical rules and transformations.
6 Conclusion
In this work, we enhance semantic representations of symbolic expressions by developing E-Gen, a novel mathematical corpus generation framework based on the e-graph data structure. E-Gen shows strong capability in generation of a scalable cluster-based corpus with high diversity in mathematical transformation. To evaluate the effectiveness of the new corpus, two training approaches based on seq2seq and contrastive learning respectively are implemented to capture equivalence relation between expressions. We demonstrate the efficacy of these embedding models for a variety of downstream tasks such as: clustering, mathematical information retrieval, mistake detection, and mathematical analogies. Crucially, we show that these semantic representations outperform prior work and GPT-4o in both quantitative and qualitative tests. This work provides an algorithmic foundation for processing symbolic mathematics and the vector-based representations can integrate easily with vector embeddings of other data modalities.
7 Limitations
This work introduces and evaluates the potential of a novel mathematical corpus generation framework to augment mathematical semantic understanding. While the results are promising, there are several areas for improvement. First, extending the dataset to include a larger variety of operators used in published datasets, such as ArXMLiv (Kohlhase et al., 2024) and ARQMath (Mansouri et al., 2022b), would make the vector embeddings more applicable to real world data.
Following that, more efficient grammar enumeration techniques, such as those based on the Earley algorithm (Earley, 1970), will allow us to more easily extend the E-Gen corpus to include expressions with more complex operators and larger arity. In this manuscript, equivalent expressions are generated by enumerating the grammar obtained from the e-graph using a simple recursive function. Additionally, incorporating variable characteristics, such as dimensionality, phase, and/or bounds, into the embeddings would be useful and an avenue yet to be explored in modern literature.
Another limitation is the scarcity of real-world datasets that focus on symbolic mathematics. While tasks like mathematical information retrieval have been explored in competitions such as NTCIR (Zanibbi et al., 2016a) and ARQMath (Mansouri et al., 2022b), these typically involve a mix of symbolic math and natural language in queries and results. Integration of the our embedding methods with natural language vectors is a research area that has yet to be explored.
Appendix A Cluster-Based Corpus
A.1 Generation Efficiency
Efficiency is a key property of any corpus generation scheme. To evaluate the E-Gen approach, the average time for e-graph saturation and expression rewrite extraction is measured. On average, e-graph saturation process takes 36.83ms, while extraction takes 1.46s per expression. The process is run on an Intel Xeon E5-2666 CPU with a 25-token length limit and a 600s time limit for each raw expression.
A.2 Cluster Example
As discussed in Section 3, the E-Gen framework enables wider coverage of operators, such as derivatives, and the application of a diverse range of mathematical rules. Table 12 shows additional examples of equivalent expression clusters. The key distinction between the E-Gen corpus and prior datasets is that E-Gen consists of clusters of equivalent expressions, rather than simple expression pairs. Each cluster contains a raw expression along with numerous equivalent rewrites. This diverse set of transformations significantly improves the performance of pretrained models in understanding the semantics of mathematical transformations.
Raw expression: $(\sin(2x+5))^{-8}$
$(\sin(5-(-2)\sin(\sin^{-1}(x))))^{-8}$
$(\csc(\pi+(-5-2x)))^{8}$
$(\cos(\pi/2-2x-5)))^{-8}$
$\left(\frac{\cot(2x+5)}{\cos(2x+5)}\right)^{8}$
$(\sin(2(x+\pi)+5))^{-8}$
Raw expression: $\sec^{-1}(7x+6)/(-7)$
$\cos^{-1}(\frac{1}{6-(-7x)})/(-7)$
$(\sec^{-1}(-7x-6)-\pi))\times\frac{1}{7}$
$(\frac{\pi}{2}-\csc^{-1}(7x+6))/(-7)$
$(-1)(\sec^{-1}(7/\sec(\cos^{-1}(x))+6))/7$
$(\cos^{-1}(\cot(\tan^{-1}(7x+6))))/(-7)$
Raw expression: $(\csc(x/7))^{-3}-(-3)$
$31/\left(\csc\left(\frac{\tan(\tan^{-1}x)}{7}\right)\right)^{3}+$ 3
$(\tan(\cot^{-1}(\sin(x/7))))^{-3}+3$
$3+\left(1/\cos(\sec^{-1}(\sin(x/7)))\right)^{3}$
$-1(-3-(\csc(x/7))^{-3})$
$(\sin(x/7))^{3}+3$
$(\sin(\csc^{-1}(\sin(7^{-1}x)))^{-3}-(-3)$
Raw expression: $\frac{d}{dx}(-4x-8)^{-4}+7x$
$(-1)\times(8-(-4x))^{-5}\times 16+7$
$-16\times(4x+8)^{-5}+7$
$7+(-16)\times\left(-1/(-4x-8)^{5}\right)$
$7-(-16)\times(-1)\times(8-(-4x))^{-5}$
$0-(-16(-4x-8)^{-5}-7)$
Raw expression: $\frac{d}{dx}(-9\sinh(-7x+2))$
$63/\operatorname{sech}(-7x+2)$
$-9/(-1/7\cosh(7x-2))$
$63\cosh(2-7x)$
$\frac{63}{2}(\exp(2-7x)+\exp(7x-2))$
$7\times 9/\operatorname{sech}(7x-2)$
Table 12: Additional examples of equivalent expressions generated with E-Gen. Expressions listed below each raw expression are all equivalent to it.
Appendix B Training Details
Both the seq2seq and contrastive learning models are trained using a transformer architecture with 8 attention heads, 6 layers, a model dimensionality of 512, a feedforward dimension of 2048, and a batch size of 256. All models are optimized using the AdamW optimizer (Loshchilov, 2017) with a learning rate of $10^{-4}$ and the CosineAnnealingWarmRestarts scheduler (Loshchilov and Hutter, 2016). Training is conducted on a cluster equipped with an AMD EPYC 7763 64-Core Processor and a 46GB L40S GPU. The training time of each model is shown in Table 13.
Model Training Time (h:min)
seq2seq 38:04
CL Mean 31:56
CL Max 29:59
Table 13: Training time of seq2seq, CL Mean, and CL Max models.
Appendix C Experiments
C.1 K-Means Clustering
Accuracy Calculation.
Since all expressions from the same cluster are labeled in the same class, the accuracy for K-Means clustering in Section 5.1 is calculated as follow.
$\textit{acc}=\frac{1}{K}\sum_{i=1}^{K}\textit{acc}_{i}$ (3)
where $\textit{acc}_{i}$ denotes the accuracy of cluster $i$, which is computed as follow.
$\textit{acc}_{i}=\frac{1}{|c_{i}|}\sum_{x_{j}\in c_{i}}\mathds{1}_{\{g(x_{j})=c_{i}\}}$ (4)
where $g(x_{j})$ denotes the cluster predicted by K-Means algorithm, and $c_{i}$ is the ground truth cluster of expression $x_{j}$.
C.2 Mistake Detection
Algorithm 1 Threshold Calculation for Mistake Detection
1:input: set of derivations $\{\boldsymbol{D}_{i}\}_{i=1}^{N}$, each containing a sequence of derivation steps $\boldsymbol{D}_{i}=\{\boldsymbol{d}_{k}\}_{k=1}^{M}$
2:output: threshold value $t$
3:$f$: transformer encoder
4:$g$: cosine similarity function
5:$x\leftarrow\texttt{[]}$]
6:for all $\boldsymbol{D}_{i}\in\{\boldsymbol{D}_{1},\ldots,\boldsymbol{D}_{N}\}$ do
7:     $\boldsymbol{Z}_{i}\leftarrow f(\boldsymbol{D}_{i})$
[…]
9:     $c\leftarrow c\setminus\{c_{\text{mistake}}\}$
10:     $c_{\min}\leftarrow\min(c)$
11:     $x\leftarrow x\cup c_{\min}$
12:end for
13:$t\leftarrow\frac{1}{|x|}\sum_{i=1}^{|x|}x_{i}$
14:Return $t$
Threshold Algorithm.
The threshold for mistake detection mentioned in Section 5.1 is computed using Algorithm 1. We first generate derivation steps based on each expression in the test set of the E-Gen corpus using SymPy, 80% of derivations are used for threshold calculation. For each derivation, we calculate the cosine similarities between consecutive steps. Then, the cosine similarities of steps with mistake are removed. We record the minimum cosine similarity of the remaining correct steps for each derivation. Finally, the average of the minimum cosine similarities is calculated as the threshold used for binary classification in the mistake detection task.
C.3 Formula Retrieval
Formula retrieval is implemented as an extra task of semantic understanding. As discussed in Section 1, prior MIR studies rely heavily on contextual cues for semantic representation (Gao et al., 2017; Krstovski and Blei, 2018) rather than really explore mathematical property of expressions, resulting in suboptimal performance, especially in scenarios with limited surrounding text, such as textbook or mathematical derivations. To evaluate the model’s potential in this scenario, we design a test simulating pure formula retrieval based on the E-Gen corpus. Specifically, given a query expression, the top-$k$ most similar expressions are retrieved based on cosine similarity in the continuous embedding space.
Model $\boldsymbol{k}$=5 $\boldsymbol{k}$=10 $\boldsymbol{k}$=15 $\boldsymbol{k}$=20
seq2seq 99.93 99.82 99.76 99.41
CL Mean 99.85 99.73 99.50 98.94
CL Max 99.78 99.52 99.32 98.82
SemEmb 73.47 60.57 51.95 45.56
Table 14: Mathematical information retrieval accuracy (%) of seq2seq, CL Mean, and CL Max, compared against prior SemEmb model at different top-$k$ values. Accuracy denotes the proportion of top-$k$ ranked candidates that are semantically equivalent to the query expression.
The test set of E-Gen corpus which contains 8,077 expressions is tested, each expression is used as a query while the remaining expressions are used as candidates. Figure 4 shows an example in which top-$4$ most similar candidates to $\cos{x}$ are selected. As Table 14 shows, both seq2seq and contrastive learning models based on E-Gen corpus can effectively retrieve the most semantically relevant expressions from the candidate pool as $k$ increases. Conversely, due to a limited number of rewrites per expression, SemEmb struggles to identify semantically equivalent expressions when $k$ exceeds 10.
Figure 4: Example of the formula retrieval with seq2seq model. The query expression is $\cos(x)$ in darker blue, and the rest are candidates. Top-$4$ expressions are retrieved. t-SNE is applied to reduce the dimensionality of the embeddings from 512 to 2.
C.4 Embedding Algebra
Table 15 and 16 present additional examples from embedding algebra task. The seq2seq model correctly predicts most answers, outperforming other models. CL Mean slightly outperforms, CL Max, SemEmb, and GPT-4o. Notably, SemEmb and GPT-4o use a similar strategy to predict “$\boldsymbol{y}_{2}$”. For instance, Tests 6 to 11 evaluate the models’ understanding of function periodicity, where the seq2seq model trained on the E-Gen cluster-based corpus accurately predicts “$\boldsymbol{y}_{2}$” in most cases. In contrast, as discussed in Section 5.2, SemEmb, similar to GPT-4o, tends to imitate the transformation between “$\boldsymbol{x}_{1}$” and “$\boldsymbol{y}_{1}$” rather than correctly applying the periodicity of the function. In Test 6, “$\boldsymbol{x}_{1}:\sin(x)$” and “$\boldsymbol{y}_{1}:\sin(x+2\pi)$” are equivalent due to the $2\pi$ period of “$\sin(x)$”. However, since “$\boldsymbol{x}_{2}:\cot(x)$” has a period of $\pi$, the correct “$\boldsymbol{y}_{2}$” should be “$\cot(x+\pi)$”. The SemEmb model just simply imitates the transformation from “$\boldsymbol{x}_{1}$” to “$\boldsymbol{y}_{1}$” by adding “$2\pi$” to “$x$” in “$\cot(x)$”, and incorrectly converts “$\cot(...)$” into “$\csc(...)$”. This comparison highlights that the models trained with E-Gen corpus shows a better understanding in mathematical rules and transformations, rather than relying on superficial "looks-like" predictions.
Figure 5: Two mistake detection conversation examples in GPT-4o prompt. The mistakes in the derivation are in red. In the left one, GPT-4o successfully detects the mistake step, but also reports a false positive mistake step. In the right one, GPT-4o fails to detect the mistake in the derivation at all.
$\boldsymbol{x}_{1}$ $\boldsymbol{y}_{1}$ $\boldsymbol{x}_{2}$ $\boldsymbol{\hat{y}}_{2}$ (seq2seq) $\boldsymbol{\hat{y}}_{2}$ (CL Mean) $\boldsymbol{y}_{gt}$
$1$ $\sin(x)$ $-\sin(-x)$ $\cos(x)$ $\cos(-x)$ $\cos(-x)$ $\cos(-x)$
$2$ $\sin(x)$ $-\sin(-x)$ $\tan(x)$ $-\tan(-x)$ $-\tan(-x)$ $-\tan(-x)$
$3$ $\cos(x)$ $\sec(x)$ $\tanh(x)$ $\coth(x)$ $\coth(x)$ $\coth(x)$
$4$ $\sin^{-1}(x)$ $\csc^{-1}(1/x)$ $\cos^{-1}(x)$ $\sec^{-1}(1/x)$ $\sec^{-1}(1/x)$ $\sec^{-1}(1/x)$
$5$ $\sinh^{-1}(x)$ $\operatorname{csch}^{-1}(1/x)$ $\tanh^{-1}(x)$ $\coth^{-1}(1/x)$ $\coth^{-1}(1/x)$ $\coth^{-1}(1/x)$
$6$ $\sin(x)$ $\sin(x+2\pi)$ $\cot(x)$ $\cot(x+\pi)$ $\cot(x+\pi)$ $\cot(x+\pi)$
$7$ $\tan(x)$ $\tan(x+\pi)$ $\csc(x)$ $\csc(x+2\pi)$ $\csc(x+2\pi)$ $\csc(x+2\pi)$
$8$ $-\cos(x)$ $\cos(x+\pi)$ $-\cot(x)$ ${\color[rgb]{1,0,0}\cot(x)\times 1}$ ${\color[rgb]{1,0,0}\cos(x)(1/\sin(x))}$ $\tan(x+\pi/2)$
$9$ $\sin(x)$ $\cos(x-\pi/2)$ $\sec(x)$ $\csc(x+\pi/2)$ $\csc(x+\pi/2)$ $\csc(x+\pi/2)$
$10$ $-\csc(x)$ $\csc(x+\pi)$ $-\sec(x)$ $\sec(x+\pi)$ ${\color[rgb]{1,0,0}\cot(x)\csc(\pi/2+x)}$ $\sec(x+\pi)$
$11$ $-\cos(x)$ $\cos(x+\pi)$ $-\tan(x)$ ${\color[rgb]{1,0,0}\tan(\pi+x)}$ ${\color[rgb]{1,0,0}\sin(-\csc^{-1}(\cot(x))+\pi)}$ $\cot(x+\pi/2)$
$12$ $x$ $\ln{x}$ $\sin(x)$ $\ln{\sin(x)}$ ${\color[rgb]{1,0,0}\sin(\ln(x))}$ $\ln{\sin(x)}$
$13$ $x$ $\ln{x}$ $\coth(x)$ $\ln{\coth(x)}$ $\ln{\coth(x)}$ $\ln{\coth(x)}$
$14$ $x$ $\ln{x}$ $\cos^{-1}(x)$ ${\color[rgb]{1,0,0}\cos^{-1}(\ln x)}$ ${\color[rgb]{1,0,0}\ln{\sec^{-1}(x)}}$ $\ln{\cos^{-1}(x)}$
$15$ $x$ $\sin^{-1}(x)$ $\cos(x)$ $\sin^{-1}(\cos(x))$ $\sin^{-1}(\cos(x))$ $\sin^{-1}(\cos(x))$
$16$ $x$ $\csc^{-1}(x)$ $\operatorname{csch}(x)$ $\csc^{-1}(\csc(x))$ ${\color[rgb]{1,0,0}\csc^{-1}(\sinh(x))}$ $\csc^{-1}(\csc(x))$
$17$ $x$ $x+1$ $\tan^{-1}(x)$ ${\color[rgb]{1,0,0}\tan^{-1}(x)\times 1}$ $\tan^{-1}(x)+1$ $\tan^{-1}(x)+1$
$18$ $x$ $x-1$ $\sin(x)$ $\sin(x)-1$ $\sin(x)-1$ $\sin(x)-1$
$19$ $x$ $x/1$ $\tan(x)$ $\tan(x)/1$ ${\color[rgb]{1,0,0}(\frac{d}{dx}\ln(x))^{1}}$ $\tan(x)/1$
$20$ $x$ $x^{3}$ $\operatorname{csch}(x)$ $\operatorname{csch}^{3}(x)$ $\operatorname{csch}^{3}(x)$ $\operatorname{csch}^{3}(x)$
$21$ $x$ $x^{3}$ $\sinh^{-1}(x)$ $\sinh^{-3}(x)$ ${\color[rgb]{1,0,0}5+(1/\tan(\cot^{-1}(x)))^{3}}$ $\sinh^{-3}(x)$
Table 15: Additional examples of embedding algebra with seq2seq and CL Mean models. Both models are trained using E-Gen corpus and this evaluation shows the difference between seq2seq and CL Mean models. $\hat{\boldsymbol{y}}_{2}$ is the prediction from the corresponding model, and $\boldsymbol{y}_{gt}$ is the ground truth. The expressions in red are wrong predictions. The seq2seq model offers superior performance over the CL Mean model.
$\boldsymbol{x}_{1}$ $\boldsymbol{y}_{1}$ $\boldsymbol{x}_{2}$ $\boldsymbol{\hat{y}}_{2}$ (CL Max) $\boldsymbol{\hat{y}}_{2}$ (SemEmb) $\boldsymbol{\hat{y}}_{2}$ (GPT-4o) $\boldsymbol{y}_{gt}$
$1$ $\sin(x)$ $-\sin(-x)$ $\cos(x)$ $\cos(-x)$ ${\color[rgb]{1,0,0}-\tan(-x)}$ $\cos(-x)$ $\cos(-x)$
$2$ $\sin(x)$ $-\sin(-x)$ $\tan(x)$ $-\tan(-x)$ $-\tan(-x)$ $-\tan(-x)$ $-\tan(-x)$
$3$ $\cos(x)$ $\sec(x)$ $\tanh(x)$ $\coth(x)$ $\coth(x)$ ${\color[rgb]{1,0,0}\operatorname{sech}(x)}$ $\coth(x)$
$4$ $\sin^{-1}(x)$ $\csc^{-1}(1/x)$ $\cos^{-1}(x)$ $\sec^{-1}(1/x)$ $\sec^{-1}(1/x)$ $\sec^{-1}(1/x)$
$5$ $\sinh^{-1}(x)$ $\operatorname{csch}^{-1}(1/x)$ $\tanh^{-1}(x)$ $\coth^{-1}(1/x)$ ${\color[rgb]{1,0,0}\coth^{-1}(x)}$ $\coth^{-1}(1/x)$
$6$ $\sin(x)$ $\sin(x+2\pi)$ $\cot(x)$ $\cot(x+\pi)$ ${\color[rgb]{1,0,0}\csc(x+2\pi)}$ $\cot(x+\pi)$ $\cot(x+\pi)$
$7$ $\tan(x)$ $\tan(x+\pi)$ $\csc(x)$ $\csc(x+2\pi)$ ${\color[rgb]{1,0,0}\cot(x+\pi)}$ $\csc(x+2\pi)$ $\csc(x+2\pi)$
$8$ $-\cos(x)$ $\cos(x+\pi)$ $-\cot(x)$ ${\color[rgb]{1,0,0}\sec(\cos^{-1}(\cot(x)))}$ ${\color[rgb]{1,0,0}\sec(x+\pi)}$ ${\color[rgb]{1,0,0}\cot(x+\pi)}$ $\tan(x+\pi/2)$
$9$ $\sin(x)$ $\cos(x-\pi/2)$ $\sec(x)$ ${\color[rgb]{1,0,0}\cos(x)}$ ${\color[rgb]{1,0,0}\sec(x-\pi/2)}$ ${\color[rgb]{1,0,0}\csc(x-\pi/2)}$ $\csc(x+\pi/2)$
$10$ $-\csc(x)$ $\csc(x+\pi)$ $-\sec(x)$ ${\color[rgb]{1,0,0}\csc(\pi/2+x)}$ $\sec(x+\pi)$ $\sec(x+\pi)$ $\sec(x+\pi)$
$11$ $-\cos(x)$ $\cos(x+\pi)$ $-\tan(x)$ ${\color[rgb]{1,0,0}\sin(\csc^{-1}(\cot(x)))}$ ${\color[rgb]{1,0,0}\tan(x+\pi)}$ ${\color[rgb]{1,0,0}\tan(x+\pi)}$ $\cot(x+\pi/2)$
$12$ $x$ $\ln{x}$ $\sin(x)$ ${\color[rgb]{1,0,0}\ln(\csc(x))}$ $\ln{\sin(x)}$ ${\color[rgb]{1,0,0}\sin^{-1}(x)}$ $\ln{\sin(x)}$
$13$ $x$ $\ln{x}$ $\coth(x)$ $\ln{\coth(x)}$ ${\color[rgb]{1,0,0}\coth{\ln(x)}}$ ${\color[rgb]{1,0,0}\coth^{-1}(x)}$ $\ln{\coth(x)}$
$14$ $x$ $\ln{x}$ $\cos^{-1}(x)$ ${\color[rgb]{1,0,0}\sec^{-1}(\ln x)}$ ${\color[rgb]{1,0,0}\cos^{-1}({\ln(x)})}$ $\ln{\cos^{-1}(x)}$ $\ln{\cos^{-1}(x)}$
$15$ $x$ $\sin^{-1}(x)$ $\cos(x)$ $\sin^{-1}(\cos(x))$ ${\color[rgb]{1,0,0}\cos(\sin^{-1}(x))}$ ${\color[rgb]{1,0,0}\cos^{-1}(x)}$ $\sin^{-1}(\cos(x))$
$16$ $x$ $\csc^{-1}(x)$ $\operatorname{csch}(x)$ ${\color[rgb]{1,0,0}\sin^{-1}(\operatorname{csch}(x))}$ ${\color[rgb]{1,0,0}\operatorname{csch}^{-1}(x)}$ $\csc^{-1}(\csc(x))$
$17$ $x$ $x+1$ $\tan^{-1}(x)$ ${\color[rgb]{1,0,0}\tan^{-1}(x)/1}$ $\tan^{-1}(x)+1$ $\tan^{-1}(x)+1$ $\tan^{-1}(x)+1$
$18$ $x$ $x-1$ $\sin(x)$ $\sin(x)-1$ ${\color[rgb]{1,0,0}\sin(x)+1}$ ${\color[rgb]{1,0,0}\sin(x-1)}$ $\sin(x)-1$
$19$ $x$ $x/1$ $\tan(x)$ ${\color[rgb]{1,0,0}\frac{d}{dx}\ln x}$ $\tan(x)/1$ $\tan(x)/1$ $\tan(x)/1$
$20$ $x$ $x^{3}$ $\operatorname{csch}(x)$ $\operatorname{csch}^{3}(x)$ $\operatorname{csch}^{3}(x)$ $\operatorname{csch}^{3}(x)$
$21$ $x$ $x^{3}$ $\sinh^{-1}(x)$ ${\color[rgb]{1,0,0}\operatorname{csch}^{-1}(x)}$ $\sinh^{-3}(x)$ $\sinh^{-3}(x)$ $\sinh^{-3}(x)$
Table 16: Additional examples of embedding algebra comparing the embedding model (CL Max) trained on the E-Gen corpus presented in this manuscript with prior work (SemEmb) and GPT-4o. $\hat{\boldsymbol{y}}_{2}$ is the prediction from the corresponding model, and $\boldsymbol{y}_{gt}$ is the ground truth. The expressions in red are wrong predictions.